OMG Inc. has 4 million shares of common stock outstanding, 3 million shares of preferred stock outstanding, and 5,000 bonds. Suppose the common shares are selling for $26 per share, the preferred shares are selling for $25 per share, and the bonds are selling for 109 percent of par.
What weight should you use for debt in the computation of OMG’s WACC? (Round your answer to 2 decimal places.)
Weight used: ___.__%
Answer #1

Total value of:

Common stock=(4*26)=$104 million

Preferred stock=(3*25)=$75 million

Bonds=5000*(1000*109%)=$5.45 million

Total value=(104+75+5.45)=184.45 million

Hence weight of debt=(5.45/184.45)

=2.95%(Approx).